OpenStudy (anonymous):

please help im so stuck Given the following functions f(x) and g(x), solve f over g (−5) and select the correct answer below: f(x) = 2x − 20 g(x) = x − 1

OpenStudy (danjs):

f(-5) / g(-5) that right?

OpenStudy (anonymous):

-5 is one of the choices but im not sure

OpenStudy (anonymous):

could you explain to me how you got that answer?

OpenStudy (anonymous):

−5 5 one sixth 30 the coices

OpenStudy (danjs):

\[\frac{ f(-5) }{ g(-5) } = \frac{ 2(-5)-20 }{ -5 -1 }\]

OpenStudy (danjs):

right , yeah it is f(-5) / g(-5), just evaluate both and put them over each other

OpenStudy (anonymous):

ok thank you can you help with few more just trying out openstudy for the first time my friends said it helps

OpenStudy (danjs):

yes, depends on who helps ya.. .. just ask a new thread

OpenStudy (anonymous):

you mean like post the question by itself?

OpenStudy (danjs):

screw it, here is fine

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

given the functions f(x)and g(x), solve (f times g)(3) and select the correct answer below f(x)4x^2+12 g(x)=x-1

OpenStudy (danjs):

f times g, or is it a small circle looking thing

OpenStudy (anonymous):

47 -47 96 -96 are choices

OpenStudy (anonymous):

its a small cirlce

OpenStudy (danjs):

ok, then it is a compound function f[g(x)]

OpenStudy (anonymous):

so how would i set up the equation

OpenStudy (danjs):

nope, it is gonna be multiplication

OpenStudy (danjs):

f(3) * g(3) =

OpenStudy (anonymous):

so i would multiply the 2 equation they gave me?

OpenStudy (anonymous):

so (4x^2+12)3 and (x-1)3?

OpenStudy (danjs):

you could do that first, or put in the value for x to start, [4(3^2) + 12] * (3-1)

OpenStudy (anonymous):

hows did you get the value of x

OpenStudy (danjs):

THe problem says [f times g ](3) , x=3

OpenStudy (danjs):

so you just evaluate both f and g at x=3 , then multiply the values together

OpenStudy (anonymous):

ooo ok

OpenStudy (anonymous):

let me try it

OpenStudy (danjs):

I have time for one more if you want

OpenStudy (danjs):

f(3) * g(3) = [4(3^2) + 12] * (3-1) = 96

OpenStudy (anonymous):

how did u get 96 i got 168

OpenStudy (danjs):

[4*9 + 12]* 2

OpenStudy (danjs):

48*2

OpenStudy (anonymous):

ok and wat a bout the part when you multiply it by 3-1

OpenStudy (danjs):

that is what happened, the 2 at the end... f(3) * g(3) = [4(3^2) + 12] * (3-1) = 96

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

i get it now i messed up omewhere but i see thay why its a 2 at the end cuz 3-1 is 2 ok can u hepl with one more?

OpenStudy (danjs):

yep

OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8

OpenStudy (anonymous):

choices are −96 0 24 48

OpenStudy (danjs):

These compound functions like that, you work from the inside most function first, then outwards carrying the values...

OpenStudy (danjs):

figure out g(6), then use that value in f(x) for x

OpenStudy (anonymous):

so x is 6?

OpenStudy (danjs):

yes, anytime you see g(#) or f(#) some number, that is what x is

OpenStudy (anonymous):

k so now i plug it in dont tell me ill tell u n u tell if its right

OpenStudy (danjs):

So what is g(6)

OpenStudy (anonymous):

is it -2?

OpenStudy (danjs):

no... First , what is g(6) ?

OpenStudy (anonymous):

i did x-8 and plugged 6 for x

OpenStudy (danjs):

right, so g(6) = 6 - 8 = -2

OpenStudy (anonymous):

ok so now f

OpenStudy (danjs):

now you use that value you just got in the f function for x

OpenStudy (anonymous):

f is 48?

OpenStudy (danjs):

f(-2)

OpenStudy (anonymous):

6times 6 plus 12

OpenStudy (danjs):

no, that is the thing with the compound functions, you have to use the value from g(6) = -2 for the value of x in f(x)

OpenStudy (danjs):

Figure the inner function out, use that value in the next function

OpenStudy (anonymous):

o ok so i plug the -2 in instead of the 6?

OpenStudy (danjs):

Yes.. f[g(6)] = f(-2)

OpenStudy (anonymous):

can you explain how you get that from f(x)=6x+12?

OpenStudy (danjs):

ok, start over from the beginning...

OpenStudy (anonymous):

ok

OpenStudy (danjs):

f [ g(6) ] What that means is instead of having the usual f(x), you have f( another function) Start with the inner function g(6) and evaluate. g(6) = -2 Now you can replace g(6) with -2 in the original problem

OpenStudy (danjs):

Left to evaluate f(-2)

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

so where do i go from here?

OpenStudy (danjs):

just left to evaluate f(-2) = 6*(-2) + 12

OpenStudy (anonymous):

o ok so now u added the 6(-2)+12 wouldnt it be 0

OpenStudy (danjs):

so it is zero i have to take off, you can fan me , ill prolly be on later

OpenStudy (anonymous):

ok how do i fan you?

OpenStudy (anonymous):

nvm n did it thanks